[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Feb 15, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Feb 15 16:11:57 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. The Derecho N-AWIPS workstation froze again with minimal SPFs loaded,  
which required the system to be powered down, resulting in a 20 minute loss  
of work time. TSB is waiting for NCO to come up with the final  
hardware/software configuration for the RHEL6 version of N-AWIPS. We cannot  
purchase the hardware upgrades until we receive the final specifications on  
what is needed. TSB will expedite the purchase as soon as NCO resolves the  
configuration issue. TSB Developers Matt Sardi and Chris Mello are  
assisting NCO as much as possible to keep this issue moving forward.

2. The AWIPS-2 patch that was applied yesterday reintroduced several  
previously encountered errors. TSB AWIPS-2 Sysadmin Chris Mello was able to  
quickly fix these problems.

3. TAFB Forecaster Nelsie Ramos noticed problems calling up and displaying  
the Offshore Waters OFFNT4 forecast  
(http://www.nhc.noaa.gov/text/MIAOFFNT4.shtml) and the NAVTEX OFFN04  
forecast (http://www.nhc.noaa.gov/text/MIAOFFN04.shtml) in the AWIPS-2  
GUI/text formatter. Old versions of the products are displayed. Chris Mello  
is working with Nelsie to solve these problems.

HSU:
1. NCO Linux Sysadmin Brian Maher backed down the NVIDIA graphics driver on  
the Tuna N-AWIPS workstation. Senior Hurricane Specialist Jack Beven  
noticed a marked improvement in the time it takes to load satellite SPFs.  
Peformance has been restored to nearly normal load times.

2. At 1428Z yesterday, Senior Hurricane Specialist Mike Brennan submitted a  
P-Surge test run on WCOSS for test storm AL832017.

3. The SSU is testing a new AWIPS-2 fix from Raytheon that should improve  
the transfer speeds of ISC (inter-site coordination) grids like the NHA SS  
domain.
Other items:
1. The security camera display on the TAFB ops wall and on the MFL ops wall  
is shrunk and needs to be expanded back to its normal size. NCO Windows  
Sysadmin Marc Laurent will fix this display.

2. The first part of the UP04 software maintenance upgrade on the Luna Cray  
was completed and baseline tests were successful. The outage lasted longer  
than anticipated, nearly 20 hours. A Tide Phase 2 storage upgrade (16 hour  
outage) and the second part of the UP04 upgrade for Luna [Cray Linux  
Environment (CLE) 5.2, 24 hour outage] will begin on Monday, Feb. 20.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170215/7173d87c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list